OpenStudy (anonymous):

Chain rule with quotient rule....? I tried to solve this problem but ended up in a numerical mess: find the derivative of :y=x/[sqrt(x^4+4)]

OpenStudy (anonymous):

quotient rule

OpenStudy (anonymous):

\[\frac{f}{g}=\frac{x}{\sqrt{x^4+4}}\]

OpenStudy (anonymous):

\[\huge (\frac{f}{g})'=\frac{\sqrt{x^4+4}-\frac{2x^4}{\sqrt{x^4+4}}}{x^4+4}\]

OpenStudy (anonymous):

not that too mess

OpenStudy (anonymous):

flawless work @mukushla, just to avoid confusion, on the quotient in the numerator, it should be 2x^3 right?

OpenStudy (anonymous):

dont forgot x f'g-fg'

OpenStudy (anonymous):

oh yes, thank you, I missed that step!.

OpenStudy (anonymous):

@rymdenbarn

OpenStudy (anonymous):

is that what u got?

OpenStudy (anonymous):

@mukushla I had slipped up towards the end but I get it now This was my original answer

OpenStudy (anonymous):

almost done